a chip is placed at room temperature of 40°c after minutes its temperature changes from 80c to 60c the proportionallity constant in case is k=0.05 ln2 find temperture of chip after 40 minute.


1
Expert's answer
2022-05-20T06:48:35-0400

Recall Newton’s law of cooling: the rate of change in temperature of an object is proportional

to the difference in temperature between the object and its surroundings,

"dT\/dt=k(T-T_s)"

"\int ( dT/(T-T_s))=\int k dt"

"ln(T-T_s)=kt+C_1"

"T=T_s+C_2e^{kt}"

T(0)=Ts+C2

C2=T(0)-Ts

"T(t)=T_s+(T(0)-T_s)e^{kt}"

"T(40)=40+(80-40)e^{0.05ln2x40}=200"